In the Bohr model of the hydrogen atom, an electron moves in a circular path around a proton. The speed of the electron is approximately 2.28 106 m/s.
(a) Find the force acting on the electron as it revolves in a circular orbit of radius 5.40 10-11 m.
magnitude N
direction

(b) Find the centripetal acceleration of the electron.
m/s2
direction
1
Expert's answer
2018-10-15T15:41:09-0400

(a) Find the force acting on the electron as it revolves in a circular orbit of radius 5.40 10-11 m.

Solution

F=m v^2/r=(9.11∙〖10〗^(-31))/(5.4∙〖10〗^(-11) ) (2.28∙〖10〗^6 )^2=8.77∙〖10〗^(-8) N.

Magnitude: 8.77∙〖10〗^(-8) N

Direction: To the nucleus.


(b) Find the centripetal acceleration of the electron.

Solution

a=v^2/r=1/(5.4∙〖10〗^(-11) ) (2.28∙〖10〗^6 )^2=9.63∙〖10〗^22 m/s^2 .

9.63∙〖10〗^22 m/s^2 .

Direction: To the nucleus.